HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

OZONE SPECTROSCOPY IN THE TERAHERTZ RANGE FROM HIGH-RESOLUTION SYNCHROTRON SOLEIL EXPERIMENTS

COMBINED WITH FAR-INFRARED MEASUREMENTS AND AB INITIO INTENSITY CALCULATIONS

\underline{V. TYUTEREV} , Tomsk State University, Tomsk, Russia; GSMA UMR CNRS 7331, Reims University, France; A. BARBE, GSMA UMR CNRS 7331, Reims University, France; \underline{L. MANCERON} , Synchrotron SOLEIL, Beamline AILES, Saint-Aubin, France; B. GROUIEZ, GSMA UMR CNRS 7331, Reims University, France ; S.

TASHKUN, Institute of Atmospheric Optics SB RAN, Tomsk, Russia ; J. BURGALAT, \underline{M. ROTGER} , GSMA UMR CNRS 7331, Reims University, France

More than half of the of the outgoing longwave radiation from the Earth to space is located in the FIR part of the electromagnetic spectrum at wavelengths greater than 15 micron. Ozone is one of the most important atmospheric species in terms of the impact on the climate changes and human health. It has a strong absorption in the MW, THz and FIR, though only a relatively restricted part of lines has been experimentally measured in these ranges, which are the most efficient ones for the synchrotron setup at the SOLEIL CNRS equipment, in terms of the brightness if the synchrotron radiation, sensitivity of measurements and S/N ratio. The experimental setup for the ozone SOLEIL project has been previously described ^{1}. In this work, the analyses of the experimental spectra recorded at four different P*L products for the main ozone isotopologue and the theoretical modeling based on an improved effective Hamiltonian for the line positions and ab initio intensity calculations^{2} ^{3} will be presented. The results will be given for the simultaneous fit of the rotational band GS-GS , the hot band ν_{2}-ν_{2} and the FIR ν_{2} band conducting to the overall weighted standard deviation of 0.8 for 13500 line positions, including all previously available MW and FIR data and the original SOLEIL measurements providing experimental accuracy of 0.00005-0.0001 cm^{−1} for the best lines. The accuracy issues and possible improvements in the SMPO, GEISA and HITRAN databases for large Ka quantum numbers will be discussed. Supports from the CNRS SOLEIL research center for the experimental setup (project N 20211156) and from the RSF project 19-12-00171-P for the theoretical work and data analyses are acknowledged.

1L.Manceron, A.Barbe, V.Tyuterev et al., Far infrared spectroscopy of the ozone molecule and its isotopomers between 50 and 800 cm^{−1}, ASA-HITRAN Conference, Reims, (2022)

2doi:10.1063/1.4973977, V.Tyuterev, R. Kochanov, S. Tashkun, Accurate ab initio dipole moment surfaces of ozone: First principle intensity predictions for rotationally resolved spectra in a large range of overtone and combination bands JCP, 146(6), art.no. 064304 (2017),

3doi:10.1016/j.jqsrt.2021.107801, V.Tyuterev, A.Barbe, S.Mikhailenko, E.Starikova, Y. Babikov, Towards the intensity consistency of the ozone bands in the infrared range: ab initio corrections to the SMPO database JQSRT, 272, art.no. 107801 (2021).

p-number: p073  Submitted on Mon, 12 Jun 2023 08:43:32 +0200